With these thoughts in mind, consider different strategies you can use to solve the following problem:
452 + 721 + 908
Solve the problem using at least two methods other than the traditional algorithm.
You may add 400+700+900=2000, then add the tens and ones 52+21+08=81, and lastly add these two numbers together 2000+81=2081.
You may add two numbers at a time: 452+721=1173+908= 2081.

Found 2 solutions by stanbon, mathtechesnba:
Answer by stanbon(75887)

What "thoughts in mind"?
Cheers,
Stan H.

Answer by mathtechesnba(1)

You may add 400+700+900=2000, then add the tens and ones 52+21+08=81, and lastly add these two numbers together 2000+81=2081.
You may add two numbers at a time: 452+721=1173+908= 2081.